OpenStudy (dls):

Inverse trigo help! [12]

OpenStudy (dls):

\[\large ~If~\cos^{-1} \lambda+\cos^{-1} u+\cos^{-1}v=3\pi,then~\lambda u+uv+v \lambda=?\]

OpenStudy (yrelhan4):

--> arccos lambada=arccos u= arccos v=pi... kyunki pi se badi value le nhi sakta arccosx ab solve krle..

OpenStudy (dls):

epic :")

OpenStudy (dls):

its not necessary that all of them are equal to pi :| it can also be done in this arrangement: \[\large \pi+\frac{\pi}{2}+\frac{3\pi}{2}\]

OpenStudy (yrelhan4):

3pi/2 arccosx ke range me nhi hai. [0, pi]

OpenStudy (dls):

hmm :|

OpenStudy (yrelhan4):

sahi keh rha hun @ParthKohli ?

OpenStudy (dls):

seems legit :-/

OpenStudy (dls):

but fishy :P

Parth (parthkohli):

@DLS Aap batayen ki mujhe trig nahi aati

OpenStudy (dls):

"NEVER CALL PARTH KOHLI FOR TRIGONOMETRY"

Parth (parthkohli):

:')

OpenStudy (anonymous):

@DLS Do you require further help or are you fine?

OpenStudy (dls):

I want to know how would we do this if we had another number on RHS lets say 6pi or 7pi

OpenStudy (yrelhan4):

No solution. ^

OpenStudy (dls):

seems legit :P

OpenStudy (dls):

another question,what If i had arc tangent instead of arc cos x function?

OpenStudy (anonymous):

No solution again.

OpenStudy (dls):

why?

OpenStudy (anonymous):

The range of artan(x) is -pi/2 < y < pi/2, i.e. -pi/2 an pi/2 are the asymptotes of arctan(x). For the sum to be 3pi would be impossible since the function gets really close to pi/2 as it goes to infinity, you could think of the values as always being less than pi/2 but very close to pi/2. Therefore, the sum arctan(u) + arctan(v) + arctan(lambada) could at most have a value of a little less than 3pi/2. Therefore, if it was equated with 3pi, there would be no solution.

OpenStudy (anonymous):

Do you understand?

OpenStudy (dls):

yes i do <3 thanks everyone!

OpenStudy (dls):

and the last case :| what if I had arctan and pi ?

OpenStudy (dls):

or arccos and pi